OpenStudy (zsahib01):

6. Write a function rule for each table. Hours Worked Pay 2 $15.00 4 $30.00 6 $45.00 8 $60.00 p = 15h p = h + 15 h = 7.50p p = 7.50h

Mehek (mehek14):

Use the slope formula to find the rate of change: \(\dfrac{y_2-y_1}{x_2-x_1}\)

Mehek (mehek14):

Pick any 2 ordered pairs, and then we'll plug them in and solve.

Mehek (mehek14):

Which 2 ordered pairs would you like to use?

OpenStudy (zsahib01):

2 and 15.00

Mehek (mehek14):

Ok that's one (2, 15) another one?

OpenStudy (zsahib01):

4 and 30.00

Mehek (mehek14):

Good Now we plug them in \(\dfrac{30-15}{4-2}=?\)

OpenStudy (zsahib01):

15/2

Mehek (mehek14):

And 15/2 is = ?

OpenStudy (zsahib01):

7.5

Mehek (mehek14):

Good so it's either C or D

Mehek (mehek14):

The pay depends of the hours so which choice do you think is the answer?

OpenStudy (zsahib01):

D?

Mehek (mehek14):

Yes correct :)

OpenStudy (zsahib01):

thank u

Mehek (mehek14):

Yw ^_^